# SceneSplat++: A Large Dataset and Comprehensive Benchmark for Language Gaussian Splatting

### v3

@@ 3 Dataset
Property SceneSplat‑7K li2025SceneSplat7k DL3DV‑10K ling2024dl3dv HoliCity zhou2021holicity Aria‑ASE avetisyan2024scenescript Crowdsourced Total
@@ 3.1 Data Collection and Processing
To ensure high-quality 3DGS scenes, we implement several quality control measures throughout the optimization pipeline. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. When depth information is available, we initialize the Gaussians at fused point cloud locations and apply depth supervision to maximize geometry quality. We use gsplat ye2025gsplat for 3DGS optimization. We filter blurry frames using the variance of the Laplacian as a sharpness metric. To efficiently compress the 3DGS scene, we employ the Markov Chain Monte Carlo strategy kheradmand20253d and add opacity and scale regularization. Once optimized, the dataset supports filtering 3DGS scenes based on PSNR and depth quality before using them as inputs for model training. We introduce the processing of each data source in the following.
SceneSplat-7K contains 7,916 curated 3DGS indoor scenes derived from ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++, Replica straub2019replica, Hypersim roberts2021hypersim, 3RScan wald2019rio, ARKitScenes baruch2021arkitscenes, and Matterport3D chang2017matterport3d.We further enrich over 6,000 of these scenes with vision–language embeddings to support open-vocabulary scene queries.
@@ 3.2 Vision-Language Embedding Collection
We follow the fusion strategy in li2025SceneSplat7k and employ a dynamic weighting mechanism that adaptively combines three distinct features: global context from the entire frame, local features with background, and masked features without background. This mechanism automatically balances contributions based on each segment’s contextual relationship—emphasizing background-inclusive features for integrated objects (e.g., keyboard with monitor) and object-specific features for isolated objects. This adaptive approach provides more nuanced semantic understanding than fixed weighting strategies.
@@ 4.1 Benchmark settings
Evaluation Protocols. We evaluate 9 methods from 3 categories on our SceneSplat-Bench benchmark. We report two key metrics in 3D: foreground mean Intersection over Union (f-mIoU) and foreground mean accuracy (f-mAcc), which address object size imbalances and reduce viewpoint dependency compared with traditional 2D-only evaluation. We unified the evaluation pipeline by first converting the language fields into a unified format $\mathcal{E}_{G}=\{E_{j}\}_{j=1}^{N}$, then computing the cosine similarity between each language feature and the queried text embeddings. Each Gaussian primitive is assigned the semantic label corresponding to the most similar text embedding. For every ground-truth 3D point used for evaluation, we identify its K-nearest Gaussian neighbors (K=25) and assign the majority-voted label as the semantic prediction for that point. We exclude floor, wall, and ceiling classes as background for indoor scenes and the sky class for outdoor scenes. To make our evaluation more complete, we additionally evaluate semantics on 2D renders, following the protocol of LERF kerr2023lerf. The 2D evaluation is conducted on subsets of the benchmark datasets with two key 2D metrics: mean Intersection over Union (mIoU) and mean accuracy (mAcc) on all classes (background and foreground classes). All runtime measurements were obtained on NVIDIA RTX A6000 GPUs, except for FMGS, whose training requires 50–80 GB of GPU memory; its runtime was therefore measured on NVIDIA H200 GPUs.
@@
LangSplat Qin2024LangSplat 0.0153 0.0741 - - 0.0057 0.0106 - -
OpenGaussian Wu2024OpenGaussians 0.1186 0.2340 - - 0.0545 0.1086 - -
Feature-3DGS zhou2024feature3DGS 0.1711 0.2512 - - 0.0541 0.0876 - -
FMGS Zuo2024FMGS 0.0916 0.2028 0.1169 0.2684 0.0641 0.1374 0.0579 0.1443
GOI QU2024GOI 0.2295 0.4638 0.2115 0.4286 0.1193 0.2297 0.1001 0.2240
@@
Gradient-Weighted 3DGS joseph2024GWfeatureBP 0.4184 0.5174 0.3573 0.4638 0.1638 0.2502 0.0601 0.1100
LUDVIG marrie2024ludvig 0.1406 0.2885 0.1337 0.2969 0.0887 0.1833 0.0602 0.1701
OccamLGS cheng2024occam 0.3559 0.4587 0.2308 0.4148 0.1936 0.2458 0.1204 0.2503
SceneSplat li2025SceneSplat7k (Pseudo Label) 0.3521 0.5210 0.3500 0.5550 0.2520 0.3038 0.2280 0.3590
@@
SceneSplat li2025SceneSplat7k 0.2968 0.5711 0.3540 0.5780 0.1387 0.4198 0.1648 0.3573
Table 3: Zero-shot 3D Semantic Segmentation Experiments on ScanNet20 dai2017scannet (20 classes) and ScanNet200 dai2017scannet (200 classes). All methods are evaluated on a 10-scene mini-validation set, with the 312-scene full validation set run only for selected methods due to runtime limitations (Tab.˜1).
@@
LangSplat Qin2024LangSplat 0.0131 0.0332 - - 0.0219 0.1257 - -
OpenGaussian Wu2024OpenGaussians 0.0695 0.1186 - - 0.1387 0.2698 - -
Feature-3DGS zhou2024feature3DGS 0.0798 0.1765 - - 0.3043 0.4732 - -
FMGS Zuo2024FMGS 0.0984 0.1892 0.1004 0.2300 0.2105 0.3664 - -
GOI QU2024GOI 0.1593 0.2790 0.1631 0.3296 0.1631 0.3296 0.1652 0.3130
@@
Gradient-Weighted 3DGS joseph2024GWfeatureBP 0.0925 0.1391 0.0900 0.1328 0.3083 0.4586 0.2762 0.4187
LUDVIG marrie2024ludvig 0.0995 0.2099 0.1099 0.2547 0.2348 0.3888 0.1960 0.3738
OccamLGS cheng2024occam 0.1580 0.2861 0.1502 0.3312 0.1803 0.3263 0.1725 0.3151
SceneSplat li2025SceneSplat7k (Pseudo Label) 0.2340 0.3709 0.2243 0.4667 0.2392 0.4368 0.2748 0.4384
@@
SceneSplat li2025SceneSplat7k 0.2263 0.4916 0.2836 0.4992 0.2732 0.5379 0.3384 0.5745
Table 4: Zero-Shot 3D Semantic Segmentation Experiments On ScanNet++ yeshwanth2023scannet++ and Matterport3D chang2017matterport3d. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
@@
   LangSplat Qin2024LangSplat    0.0193    0.0700    0.0133    0.0272    0.0111    0.0272
   OpenGaussian Wu2024OpenGaussians    0.1958    0.3294    0.0740    0.1273    0.1290    0.2133
   Feature-3DGS zhou2024feature3DGS    0.5054    0.6474    0.1867    0.2798    0.1605    0.2368
   FMGS Zuo2024FMGS    0.1413    0.2629    0.0848    0.1825    0.1134    0.2244
   GOI QU2024GOI    0.2866    0.4536    0.1134    0.1806    0.1635    0.2767
@@
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.4533    0.5785    0.1639    0.2367    0.1210    0.1760
   LUDVIG marrie2024ludvig    0.1673    0.3058    0.0958    0.1828    0.1239    0.2588
   OccamLGS cheng2024occam    0.2937    0.4645    0.1300    0.2146    0.1670    0.2866
@@
   SceneSplat li2025SceneSplat7k    0.4458    0.6727    0.2599    0.4049    0.3165    0.5094
Table 5: Zero-Shot 2D Semantic Segmentation Experiments on ScanNet20 dai2017scannet (20 classes), ScanNet200 dai2017scannet (200 classes) and ScanNet++ yeshwanth2023scannet++. All methods are evaluated on a 10-scene validation set.
@@
   LangSplat Qin2024LangSplat    0.0918    0.1950    -    -
   OpenGaussian Wu2024OpenGaussians    0.1853    0.2567    -    -
   Feature‑3DGS zhou2024feature3DGS    0.2479    0.5646    -    -
   FMGS Zuo2024FMGS    0.1774    0.2513    -    -
   GOI QU2024GOI    0.1126    0.3773    0.1587    0.3772
@@
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.2484    0.4808    0.2762    0.4187
   LUDVIG marrie2024ludvig    0.1593    0.2296    0.1843    0.2775
   OccamLGS cheng2024occam    0.2255    0.4754    0.2234    0.5056
   SceneSplat li2025SceneSplat7k (Pseudo Label)    0.2464    0.4780    0.2522    0.5159
@@
   SceneSplat li2025SceneSplat7k    0.3081    0.5647    0.2880    0.6045
Table 6: Zero-Shot 3D Outdoor Segmentation On HoliCity zhou2021holicity Dataset. All methods are evaluated on mini-validation sets,with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
@@
   LangSplat Qin2024LangSplat    0.0663    0.0415    0.0655    0.0337
   OpenGaussian Wu2024OpenGaussians    0.4565    0.3483    0.5138    0.3394
   Feature-3DGS zhou2024feature3DGS    0.5861    0.4759    0.5046    0.2783
   FMGS Zuo2024FMGS    0.6082    0.4639    0.5494    0.3303
   GOI QU2024GOI    0.5269    0.4414    0.5107    0.3568
@@
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.5486    0.4669    0.4322    0.2661
   LUDVIG marrie2024ludvig    0.6750    0.5481    0.4801    0.2875
   OccamLGS cheng2024occam    0.2074    0.1599    0.1896    0.1182
@@
   SceneSplat li2025SceneSplat7k    0.7701    0.6928    0.7339    0.5464
Table 7: 3D Object Localization Experiments on ScanNet dai2017scannet and ScanNet++ yeshwanth2023scannet++. Accuracy is reported with bounding box–based and segmentation–based evaluation.
@@
Table 8: Impact of Training-Data Scaling on Indoor Benchmarks and Cross-Domain Generalization to HoliCity zhou2021holicity. Results show that more training data consistently improves indoor performance. Furthermore, models trained on indoor data only surprisingly transfer to outdoor scenes.
@@ 4.2 Key Insights
The primary experimental results from our SceneSplat-Bench benchmark across all selected methods are presented in Tabs.˜3, 4 and 5 for indoor scenes and in Tab.˜6 for outdoor scenes. The corresponding runtime statistics for individual methods can be found in Tab.˜1. SceneSplat – representing the generalizable paradigm – stands out as the clear winner in terms of both performance and efficiency (see Fig.˜3 for qualitative examples). Interestingly, in 75% of the experiments, it even outperforms its own pseudo labels used for pretraining (SceneSplat (Pseudo Label)). This result highlights its ability to generalize by leveraging large-scale data priors and learning to predict meaningful vision-language features, effectively mitigating the noise inherent in weak supervision. In addition, Tab.˜5 shows that the performance on the 2D metric is aligned with the zero-shot performance measured with the 3D metric, with a generalizable method leading the performance. As illustrated in Fig.˜4, visualizations of text query results using enhanced NeRFView tools nerfstudio; yang2024spectrally; li2025SceneSplat7k further demonstrate that the generalizable method outperforms the others. Moreover, unlike other methods that rely on extensive feature extraction, the generalizable approach requires only 3DGS as input at inference time and predicts per-primitive vision-language features in a single forward pass.
@@
To further diversify the open-vocabulary evaluation, we include a 3D object localization task that assesses how well each method retrieves objects in 3D space using its constructed 3DGS field. Following the protocol of LangSplat Qin2024LangSplat, we directly select the point with the highest relevance score for each query. Our evaluation is conducted on 2,895 objects from ScanNet and ScanNet++ validation scenes, which is over 12× more objects compared to LERF kerr2023lerf, offering a broader and more realistic assessment. This evaluation is presented in Tab.˜7 and shows similar conclusion as previously discussed evaluations. The generalizable paradigm shows dominant performance, followed by the per-scene optimization-free methods, which outperform per-scene optimization method group.

### v4

@@ 3 Dataset
Property SceneSplat‑7K Li et al. (2025) DL3DV‑10K Ling et al. (2024) HoliCity Zhou et al. (2021) Aria‑ASE Avetisyan et al. (2024) Crowdsourced Total
@@ 3.1 Data Collection and Processing
To ensure high-quality 3DGS scenes, we implement several quality control measures throughout the optimization pipeline. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. When depth information is available, we initialize the Gaussians at fused point cloud locations and apply depth supervision to maximize geometry quality. We use gsplat Ye et al. (2025) for 3DGS optimization. We filter blurry frames using the variance of the Laplacian as a sharpness metric. To efficiently compress the 3DGS scene, we employ the Markov Chain Monte Carlo strategy Kheradmand et al. (2025) and add opacity and scale regularization. Once optimized, the dataset supports filtering 3DGS scenes based on PSNR and depth quality before using them as inputs for model training. We introduce the processing of each data source in the following.
SceneSplat-7K contains 7,916 curated 3DGS indoor scenes derived from ScanNet Dai et al. (2017), ScanNet++ Yeshwanth et al. (2023), Replica Straub et al. (2019), Hypersim Roberts et al. (2021), 3RScan Wald et al. (2019), ARKitScenes Baruch et al. (2021), and Matterport3D Chang et al. (2017).We further enrich over 6,000 of these scenes with vision–language embeddings to support open-vocabulary scene queries.
@@ 3.2 Vision-Language Embedding Collection
We follow the fusion strategy in Li et al. (2025) and employ a dynamic weighting mechanism that adaptively combines three distinct features: global context from the entire frame, local features with background, and masked features without background. This mechanism automatically balances contributions based on each segment’s contextual relationship—emphasizing background-inclusive features for integrated objects (e.g., keyboard with monitor) and object-specific features for isolated objects. This adaptive approach provides more nuanced semantic understanding than fixed weighting strategies.
@@ 4.1 Benchmark settings
Evaluation Protocols. We evaluate 9 methods from 3 categories on our SceneSplat-Bench benchmark. We report two key metrics in 3D: foreground mean Intersection over Union (f-mIoU) and foreground mean accuracy (f-mAcc), which address object size imbalances and reduce viewpoint dependency compared with traditional 2D-only evaluation. We unified the evaluation pipeline by first converting the language fields into a unified format $\mathcal{E}_{G}=\{E_{j}\}_{j=1}^{N}$, then computing the cosine similarity between each language feature and the queried text embeddings. Each Gaussian primitive is assigned the semantic label corresponding to the most similar text embedding. For every ground-truth 3D point used for evaluation, we identify its K-nearest Gaussian neighbors (K=25) and assign the majority-voted label as the semantic prediction for that point. We exclude floor, wall, and ceiling classes as background for indoor scenes and the sky class for outdoor scenes. To make our evaluation more complete, we additionally evaluate semantics on 2D renders, following the protocol of LERF Kerr et al. (2023). The 2D evaluation is conducted on subsets of the benchmark datasets with two key 2D metrics: mean Intersection over Union (mIoU) and mean accuracy (mAcc) on all classes (background and foreground classes). All runtime measurements were obtained on NVIDIA RTX A6000 GPUs, except for FMGS, whose training requires 50–80 GB of GPU memory; its runtime was therefore measured on NVIDIA H200 GPUs.
@@
LangSplat Qin et al. (2024) 0.0153 0.0741 - - 0.0057 0.0106 - -
OpenGaussian Wu et al. (2024b) 0.1186 0.2340 - - 0.0545 0.1086 - -
Feature-3DGS Zhou et al. (2024) 0.1711 0.2512 - - 0.0541 0.0876 - -
FMGS Zuo et al. (2024) 0.0916 0.2028 0.1169 0.2684 0.0641 0.1374 0.0579 0.1443
GOI Qu et al. (2024) 0.2295 0.4638 0.2115 0.4286 0.1193 0.2297 0.1001 0.2240
@@
Gradient-Weighted 3DGS Joseph et al. (2024) 0.4184 0.5174 0.3573 0.4638 0.1638 0.2502 0.0601 0.1100
LUDVIG Marrie et al. (2024) 0.1406 0.2885 0.1337 0.2969 0.0887 0.1833 0.0602 0.1701
OccamLGS Cheng et al. (2024) 0.3559 0.4587 0.2308 0.4148 0.1936 0.2458 0.1204 0.2503
SceneSplat Li et al. (2025) (Pseudo Label) 0.3521 0.5210 0.3500 0.5550 0.2520 0.3038 0.2280 0.3590
@@
SceneSplat Li et al. (2025) 0.2968 0.5711 0.3540 0.5780 0.1387 0.4198 0.1648 0.3573
Table 3: Zero-shot 3D Semantic Segmentation Experiments on ScanNet20 Dai et al. (2017) (20 classes) and ScanNet200 Dai et al. (2017) (200 classes). All methods are evaluated on a 10-scene mini-validation set, with the 312-scene full validation set run only for selected methods due to runtime limitations (Tab.˜1).
@@
LangSplat Qin et al. (2024) 0.0131 0.0332 - - 0.0219 0.1257 - -
OpenGaussian Wu et al. (2024b) 0.0695 0.1186 - - 0.1387 0.2698 - -
Feature-3DGS Zhou et al. (2024) 0.0798 0.1765 - - 0.3043 0.4732 - -
FMGS Zuo et al. (2024) 0.0984 0.1892 0.1004 0.2300 0.2105 0.3664 - -
GOI Qu et al. (2024) 0.1593 0.2790 0.1631 0.3296 0.1631 0.3296 0.1652 0.3130
@@
Gradient-Weighted 3DGS Joseph et al. (2024) 0.0925 0.1391 0.0900 0.1328 0.3083 0.4586 0.2762 0.4187
LUDVIG Marrie et al. (2024) 0.0995 0.2099 0.1099 0.2547 0.2348 0.3888 0.1960 0.3738
OccamLGS Cheng et al. (2024) 0.1580 0.2861 0.1502 0.3312 0.1803 0.3263 0.1725 0.3151
SceneSplat Li et al. (2025) (Pseudo Label) 0.2340 0.3709 0.2243 0.4667 0.2392 0.4368 0.2748 0.4384
@@
SceneSplat Li et al. (2025) 0.2263 0.4916 0.2836 0.4992 0.2732 0.5379 0.3384 0.5745
Table 4: Zero-Shot 3D Semantic Segmentation Experiments On ScanNet++ Yeshwanth et al. (2023) and Matterport3D Chang et al. (2017). All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
@@
   LangSplat Qin et al. (2024)    0.0193    0.0700    0.0133    0.0272    0.0111    0.0272
   OpenGaussian Wu et al. (2024b)    0.1958    0.3294    0.0740    0.1273    0.1290    0.2133
   Feature-3DGS Zhou et al. (2024)    0.5054    0.6474    0.1867    0.2798    0.1605    0.2368
   FMGS Zuo et al. (2024)    0.1413    0.2629    0.0848    0.1825    0.1134    0.2244
   GOI Qu et al. (2024)    0.2866    0.4536    0.1134    0.1806    0.1635    0.2767
@@
   Gradient-Weighted 3DGS Joseph et al. (2024)    0.4533    0.5785    0.1639    0.2367    0.1210    0.1760
   LUDVIG Marrie et al. (2024)    0.1673    0.3058    0.0958    0.1828    0.1239    0.2588
   OccamLGS Cheng et al. (2024)    0.2937    0.4645    0.1300    0.2146    0.1670    0.2866
@@
   SceneSplat Li et al. (2025)    0.4458    0.6727    0.2599    0.4049    0.3165    0.5094
Table 5: Zero-Shot 2D Semantic Segmentation Experiments on ScanNet20 Dai et al. (2017) (20 classes), ScanNet200 Dai et al. (2017) (200 classes) and ScanNet++ Yeshwanth et al. (2023). All methods are evaluated on a 10-scene validation set.
@@
   LangSplat Qin et al. (2024)    0.0918    0.1950    -    -
   OpenGaussian Wu et al. (2024b)    0.1853    0.2567    -    -
   Feature‑3DGS Zhou et al. (2024)    0.2479    0.5646    -    -
   FMGS Zuo et al. (2024)    0.1774    0.2513    -    -
   GOI Qu et al. (2024)    0.1126    0.3773    0.1587    0.3772
@@
   Gradient-Weighted 3DGS Joseph et al. (2024)    0.2484    0.4808    0.2762    0.4187
   LUDVIG Marrie et al. (2024)    0.1593    0.2296    0.1843    0.2775
   OccamLGS Cheng et al. (2024)    0.2255    0.4754    0.2234    0.5056
   SceneSplat Li et al. (2025) (Pseudo Label)    0.2464    0.4780    0.2522    0.5159
@@
   SceneSplat Li et al. (2025)    0.3081    0.5647    0.2880    0.6045
Table 6: Zero-Shot 3D Outdoor Segmentation On HoliCity Zhou et al. (2021) Dataset. All methods are evaluated on mini-validation sets,with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
@@
   LangSplat Qin et al. (2024)    0.0663    0.0415    0.0655    0.0337
   OpenGaussian Wu et al. (2024b)    0.4565    0.3483    0.5138    0.3394
   Feature-3DGS Zhou et al. (2024)    0.5861    0.4759    0.5046    0.2783
   FMGS Zuo et al. (2024)    0.6082    0.4639    0.5494    0.3303
   GOI Qu et al. (2024)    0.5269    0.4414    0.5107    0.3568
@@
   Gradient-Weighted 3DGS Joseph et al. (2024)    0.5486    0.4669    0.4322    0.2661
   LUDVIG Marrie et al. (2024)    0.6750    0.5481    0.4801    0.2875
   OccamLGS Cheng et al. (2024)    0.2074    0.1599    0.1896    0.1182
@@
   SceneSplat Li et al. (2025)    0.7701    0.6928    0.7339    0.5464
Table 7: 3D Object Localization Experiments on ScanNet Dai et al. (2017) and ScanNet++ Yeshwanth et al. (2023). Accuracy is reported with bounding box–based and segmentation–based evaluation.
@@
Table 8: Impact of Training-Data Scaling on Indoor Benchmarks and Cross-Domain Generalization to HoliCity Zhou et al. (2021). Results show that more training data consistently improves indoor performance. Furthermore, models trained on indoor data only surprisingly transfer to outdoor scenes.
@@ 4.2 Key Insights
The primary experimental results from our SceneSplat-Bench benchmark across all selected methods are presented in Tabs.˜3, 4 and 5 for indoor scenes and in Tab.˜6 for outdoor scenes. The corresponding runtime statistics for individual methods can be found in Tab.˜1. SceneSplat – representing the generalizable paradigm – stands out as the clear winner in terms of both performance and efficiency (see Fig.˜3 for qualitative examples). Interestingly, in 75% of the experiments, it even outperforms its own pseudo labels used for pretraining (SceneSplat (Pseudo Label)). This result highlights its ability to generalize by leveraging large-scale data priors and learning to predict meaningful vision-language features, effectively mitigating the noise inherent in weak supervision. In addition, Tab.˜5 shows that the performance on the 2D metric is aligned with the zero-shot performance measured with the 3D metric, with a generalizable method leading the performance. As illustrated in Fig.˜4, visualizations of text query results using enhanced NeRFView tools Tancik et al. (2023); Yang et al. (2024b); Li et al. (2025) further demonstrate that the generalizable method outperforms the others. Moreover, unlike other methods that rely on extensive feature extraction, the generalizable approach requires only 3DGS as input at inference time and predicts per-primitive vision-language features in a single forward pass.
@@
To further diversify the open-vocabulary evaluation, we include a 3D object localization task that assesses how well each method retrieves objects in 3D space using its constructed 3DGS field. Following the protocol of LangSplat Qin et al. (2024), we directly select the point with the highest relevance score for each query. Our evaluation is conducted on 2,895 objects from ScanNet and ScanNet++ validation scenes, which is over 12× more objects compared to LERF Kerr et al. (2023), offering a broader and more realistic assessment. This evaluation is presented in Tab.˜7 and shows similar conclusion as previously discussed evaluations. The generalizable paradigm shows dominant performance, followed by the per-scene optimization-free methods, which outperform per-scene optimization method group.